Provide process and formulas in excel

QUESTION 3

Bramble Corporation is a small wholesaler of gourmet food products. Data regarding the store's operations follow:

  • Sales are budgeted at $210,000 for November, $190,000 for December, and $180,000 for January.
  • Collections are expected to be 50% in the month of sale and 50% in the month following the sale.
  • The cost of goods sold is 55% of sales.
  • The company would like maintain ending merchandise inventories equal to 45% of the next month's cost of goods sold. Payment for merchandise is made in the month following the purchase.
  • Other monthly expenses to be paid in cash are $22,700.
  • Monthly depreciation is $13,700.
  • Ignore taxes.

Balance Sheet

October 31

Assets

Cash

$

22,500

Accounts receivable

72,500

Merchandise inventory

51,975

Property, plant and equipment, net of $574,500    accumulated depreciation

1,096,500

Total assets

$

1,243,475

Liabilities and Stockholders' Equity

Accounts payable

$

256,500

Common stock

822,500

Retained earnings

164,475

Total liabilities and stockholders' equity

$

1,243,475

Required:

Prepare the dircet material budget for December merchandise purchases:
Answer #1

- m December January 190000 180000 104500 99000 in Sales Cost of goods sold (55% of sales) Add: Desired ending merchandise in

Excel formulas:

Sales Cost of goods sold (55% of sales) December 190000 =C2*0.55 January 180000 =D2*0.55 Add: Desired ending merchandise inve